OpenStudy (anonymous):

If cot A + cosec A = 1.5 , then show that cos A= 5/13..

OpenStudy (anonymous):

reply ASAP please

OpenStudy (anonymous):

\[1+\cot ^{2}A=\csc ^{2}A\]\[\csc ^{2}A-\cot ^{2}A=1\]\[\left( \csc A+\cot A \right)\left( \csc A-\cot A \right)=1\]\[\frac{3}{2}\left( \csc A-\cot A \right)=1\]\[\csc A-\cot A=\frac{2}{3}\]\[\left( \csc A-\cot A \right)+\left( \csc A+\cot A \right)=\frac{2}{3}+\frac{3}{2}\]\[\csc A=\frac{13}{12}\]\[\sin A=\frac{12}{13}\]\[\cos A=\sqrt{1-\sin ^{2}A}=\frac{5}{13}\]

OpenStudy (anonymous):

thanks

OpenStudy (experimentx):

excellent way ... i never had seen one that way.

OpenStudy (anonymous):

I learned that way from my previous math competition problems :)

OpenStudy (anonymous):

1+cot2A=csc2A csc2A−cot2A=1 (cscA+cotA)(cscA−cotA)=1 32(cscA−cotA)=1 cscA−cotA=23 (cscA−cotA)+(cscA+cotA)=23+32 cscA=1312 sinA=1213 cosA=1−sin2A−−−−−−−−√=513

OpenStudy (anonymous):

I like @exraven's approach. +1

OpenStudy (anonymous):

However I would have done it by solving the quadratic.

OpenStudy (anonymous):

\[ 13x^2 +8x-5=0 \text{ where } x= \cos x \]

OpenStudy (experimentx):

same here ... the classic way.

OpenStudy (anonymous):

Unless it's saves time (drastically) classic rules for me! :)